calculate the molarity of an NH3 solution made up of 30.0 g of NH3 IN102ml of soln.




1
Expert's answer
2022-01-14T03:15:01-0500

Molar mass of NH3= 17.01

= 30/17.01= 1.8

= 1.8/0.102=17.65mol/L